# Establishing Linear Surrogate Regret Bounds for Convex Smooth Losses via Convolutional Fenchel–Young Losses

### v2

@@ 5 Discussions
To see this, let us denote the minimizer of (19) by $\bm{\nu}^{*}$. Then (19) is an equivalent optimization problem to (10) because there exists $\bm{\pi}\in\Pi(\bm{\theta})$ such that $\bm{\nu}^{*}=\mathcal{L}^{\bm{\rho}}\bm{\pi}$. Eventually $\nabla\Omega^{*}(\bm{\theta}+\bm{\nu}^{*})$ serves as an alternative to the gradient formula (13). Thus we can reduce the dimensionality of the optimization problem. For example, in multilabel classification (Section 2.1), $V=[0,1]^{\varrho}$ has a logarithmically smaller optimization dimensionality than $N=2^{\varrho}$. We can solve (19) with this box constraint efficiently by using the standard L-BFGS solver [30].
@@
$\mathbb{E}_{\tilde{\varphi}(\bm{\theta})}[\mathtt{Regret}_{\ell}(\tilde{\varphi}(\bm{\theta}),\bm{\eta})]=\sum_{t=1}^{N}\pi_{t}(\bm{\theta})\,\mathtt{Regret}_{\ell}(t,\bm{\eta})\leq\mathtt{Regret}_{L_{\Omega_{T}}}\!(\bm{\theta},\bm{\eta})\quad\forall(\bm{\theta},\bm{\eta})\in\mathbb{R}^{\varrho}\times\Delta^{K},$
@@
where the expectation is taken over the randomness of $\tilde{\varphi}$. Although we adopt the expected target regret differently from Theorem 13, the constant of regret bounds is strikingly improved from $N$ to $1$, which is dimension-free. Sakaue et al. [59] observes a similar regret improvement by a randomized link for online structured prediction with Fenchel–Young losses.
Strong convexity.
Since the convolutional Fenchel–Young loss possesses both smoothness and a link with a linear surrogate regret bound, it cannot be strongly convex. This can be obtained as an immediate corollary of Frongillo and Waggoner [25, Theorem 4], which proves the impossibility of the combination of smoothness, strong convexity, and a linear surrogate regret bound.333 Strictly speaking, Frongillo and Waggoner [25] assumes a target loss is non-redundant and a surrogate loss is smooth and locally strongly convex at a decision boundary.
This incompatibility can also be observed from the construction of the convolutional Fenchel–Young losses: $\Omega_{T}$ is non-smooth under non-redundant $\ell$ due to the piecewise linear $\min_{t}\left\langle\bm{p},\bm{\ell}^{\bm{\rho}}(t)\right\rangle$ and thus its conjugate is not strongly convex due to the duality between smoothness and strong convexity [28]. To put it differently, $L_{\Omega_{T}}$ cannot be strongly convex.
Acknoledgement
YC is supported by Google PhD Fellowship program. HB is supported by JST PRESTO (Grant No. JPMJPR24K6), Japan.
Appendix A Deferred Proofs in Section 3
@@
For a clear representation in this section, we denote by $T(\bm{p})$ the polyhedral convex function $-\min_{t\in\widehat{\mathcal{Y}}}\left\langle\bm{p},\bm{\ell}^{\bm{\rho}}(t)\right\rangle$, which is the negative of the affinely transformed target Bayes risk in (9). Here, $T$ is proper convex and l.s.c., and we have $\Omega_{T}=\Omega+T$.
A.1 Proof of Lemma 8
Proof.  Since $\Omega$ and $T$ are proper convex and l.s.c., so are $\Omega_{T}$ and thus $\Omega_{T}^{*}$. Then we prove (10) using the infimal convolution. First, by noting that $T$ is nothing else but the support function of the closed convex set $\mathrm{conv}(\{-\bm{\ell}^{\bm{\rho}}(t)\}_{t\in\widehat{\mathcal{Y}}})$, we can express $T^{*}$ as follows:
@@ A.1 Proof of Lemma 8
$T^{*}(\bm{\theta})=\sup_{\bm{p}\in\mathbb{R}^{\varrho}}\left[\langle\bm{\theta},\bm{p}\rangle-\max_{t\in\widehat{\mathcal{Y}}}\left\langle\bm{p},-\bm{\ell}^{\bm{\rho}}(t)\right\rangle\right]=\mathbb{I}_{\mathrm{conv}\left(\{-\bm{\ell}^{\bm{\rho}}(t)\}_{t\in\widehat{\mathcal{Y}}}\right)}(\bm{\theta}),$
where we use the conjugacy relationship between a support function and indicator function of a closed convex set $\mathrm{conv}(\{-\bm{\ell}^{\bm{\rho}}(t)\}_{t\in\widehat{\mathcal{Y}}})$ [57, Section 13]. According to Condition 1 and the definition of $T$, we see that both $\Omega$ and $T$ are proper convex and $T$ is continuous on $\mathrm{dom}(\Omega)$. Then we use the infimal convolution [13] to derive the conjugate of $\Omega_{T}$:
@@
Next we show that the infimum is indeed achieved by some $\bm{\pi}\in\Delta^{N}$. Note that for any $\bm{\theta}\in\mathbb{R}^{\varrho}$, the set $\Theta\coloneqq\{\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi}:\bm{\pi}\in\Delta^{N}\}$ is compact and non-empty, and that $\Omega^{*}$ is l.s.c. on $\Theta$ since $\mathrm{dom}({\Omega^{*}})=\mathbb{R}^{\varrho}$. Then the infimum of $\Omega^{*}(\bm{\theta})$ is achieved by some $\bm{\theta}^{\prime\prime}\in\Theta$, and there must exist $\bm{\pi}\in\Delta^{N}$ such that $\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi}=\bm{\theta}^{\prime\prime}$ by the definition of $\Theta$, which complete the proof of the existence of the minimizer.
Finally we see that for any $\bm{\theta}\in\mathbb{R}^{\varrho}$, there exists $\bm{\pi}\in\Delta^{N}$ such that $\Omega_{T}^{*}(\bm{\theta})=\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi})$, which is smaller than +$\infty$+ ∞ since $\mathrm{dom}(\Omega^{*})=\mathbb{R}^{\varrho}$ and $\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi}\in\mathbb{R}^{\varrho}$. This indicates that $\mathrm{dom}(\Omega_{T}^{*})=\mathbb{R}^{\varrho}$ because $\bm{\theta}$ is chosen arbitrarily.
A.2 Proof of Theorem 11
@@
When $\Omega$ is strictly convex, $\Omega_{T}$ is also strictly convex since $T$ is convex. Since the strict convexity holds on $\mathrm{dom}(\Omega_{T})$, $\Omega_{T}$ is further essentially strictly convex, that is, strictly convex on every convex subset of $\mathrm{dom}(\partial\Omega_{T})$ [57, p253]. According to Rockafellar [57, Theorem 26.3], $\Omega_{T}^{*}$ is essentially smooth, that is, differentiable throughout non-empty $\mathrm{int}(\mathrm{dom}(\Omega_{T}^{*}))$ with $\|\nabla\Omega_{T}^{*}(\bm{\theta})\|$ diverging to +$\infty$+ ∞ when $\bm{\theta}$ approaches a boundary point of $\mathrm{int}(\mathrm{dom}(\Omega_{T}^{*}))$, which immediately indicates the differentiability on $\mathbb{R}^{\varrho}$.
@@
A.3 Proof of Lemma 12
Proof.  According to Condition 1, the strict convexity of $\Omega$, and the proof of Theorem 11, both $\Omega_{T}^{*}$ and $\Omega^{*}$ are differentiable on $\mathbb{R}^{\varrho}$. In addition, we have $\mathrm{dom}(\Omega)=\mathrm{dom}(\Omega_{T})$ according to (8).
@@ A.3 Proof of Lemma 12
$\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi})$ =(A)${=}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi})^{\top}\nabla\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi})-\Omega(\nabla\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi}))$
@@
(A)${\leq}\bm{\theta}^{\top}\nabla\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi})+\left\langle\nabla\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi}),\mathcal{L}^{\bm{\rho}}\bm{e}_{t}\right\rangle-\Omega(\nabla\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi}))$
@@
=(B)${=}\bm{\theta}^{\top}\nabla\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi})+\underbrace{\min_{t^{\prime}\in\widehat{\mathcal{Y}}}\langle\nabla\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi}),\bm{\ell}^{\bm{\rho}}(t^{\prime})\rangle}_{=-T\left(\nabla\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi})\right)}-\Omega(\nabla\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi}))$
=(C)${=}\bm{\theta}^{\top}\nabla\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi})-\Omega_{T}(\nabla\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi}))$
(D)${\leq}\sup_{\bm{p}\in\mathrm{dom}(\Omega_{T})}\bm{\theta}^{\top}\bm{p}-\Omega_{T}(\bm{p})$
@@
where (A) owes to the optimality of $\bm{\pi}\in\Pi(\bm{\theta})$, (B) holds by the definition of $t$, and (C) holds by the definition of the convolutional negentropy $\Omega_{T}$ (8). Since $\bm{\pi}\in\Pi(\bm{\theta})$, we have $\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi})=\Omega_{T}^{*}(\bm{\theta})$ according to Lemma 8. Thus the above inequality is indeed an identity. In particular, (D) becomes an identity, which implies that the supremum of $\sup_{\bm{p}\in\mathrm{dom}(\Omega_{T})}\bm{\theta}^{\top}\bm{p}-\Omega_{T}(\bm{p})$ is achieved at $\nabla\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi})$. Since the supremum is also achieved at $\nabla\Omega_{T}^{*}(\bm{\theta})$ and the maximizer is unique according to Rockafellar and Wets [58, Proposition 11.3] and the differentiability of $\Omega^{*}_{T}$, we have
@@
Since $\bm{\pi}\in\Pi(\bm{\theta})$ and $\bm{\theta}\in\mathbb{R}^{\varrho}$ are chosen arbitrarily, this concludes the proof.
A.4 Proof of Lemma 14

### v3

@@ 5 Discussion
To see this, let us denote the minimizer of (19) by $\bm{\nu}^{*}$. Then (19) is an equivalent optimization problem to (10) because there exists $\bm{\pi}\in\Pi(\bm{\theta})$ such that $\bm{\nu}^{*}=\mathcal{L}^{\bm{\rho}}\bm{\pi}$. Eventually $\nabla\Omega^{*}(\bm{\theta}+\bm{\nu}^{*})$ serves as an alternative to the gradient formula (13). Thus we can reduce the dimensionality of the optimization problem. For example, in multilabel classification (Section 2.1), $V=[0,1]^{d}$ has a logarithmically smaller optimization dimensionality than $N=2^{d}$. We can solve (19) with this box constraint efficiently by using the standard L-BFGS solver [Liu1989].
@@
$\mathbb{E}_{\tilde{\varphi}(\bm{\theta})}[\mathtt{Regret}_{\ell}(\tilde{\varphi}(\bm{\theta}),\bm{\eta})]=\sum_{t=1}^{N}\pi_{t}(\bm{\theta})\,\mathtt{Regret}_{\ell}(t,\bm{\eta})\leq\mathtt{Regret}_{L_{\Omega_{T}}}\!(\bm{\theta},\bm{\eta})\quad\forall(\bm{\theta},\bm{\eta})\in\mathbb{R}^{d}\times\Delta^{K},$
@@
where the expectation is taken over the randomness of $\tilde{\varphi}$. Although we adopt the expected target regret differently from Theorem 13, the constant of regret bounds is strikingly improved from $N$ to $1$, which is dimension-free. sakaue2024online observes a similar regret improvement by a randomized link for online structured prediction with Fenchel–Young losses.
6 Conclusion
In this work, we construct convex and smooth surrogate losses with linear surrogate regret bounds by leveraging Fenchel–Young losses and infimal convolution. Our results demonstrate that convexity, smoothness, and linear surrogate regret are compatible for arbitrary discrete prediction problems. Moreover, our loss naturally admits consistent probability estimator, bridging the gap between linear regret and estimation. We illustrate the broad applicability of our approach through examples in multiclass classification, classification with rejection, and multilabel ranking. Overall, this study highlights the utility of convex analysis as a principled tool for designing surrogate losses.
Acknowledgement
We thank Shinsaku Sakaue for his valuable insights on the target-loss decomposition and for his careful review, and we also thank the anonymous reviewers for their attentive reading of the manuscript and their many thoughtful comments and suggestions. This research is supported by the Ministry of Education, Singapore, under its MOE AcRF Tier 2 Award MOE-T2EP20223-0003. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of the Ministry of Education, Singapore. YC is also supported by Google PhD Fellowship program. HB is supported by JST PRESTO (Grant No. JPMJPR24K6), Japan. Lei Feng is supported by the Big Data Computing Center of Southeast University.
Appendix A Additional Discussions on Related Losses
Fast rate losses.
Let us briefly discuss the target risk estimation error rates induced by different surrogate losses. Suppose that a surrogate risk can be estimated with the estimation error upper bound $\epsilon_{\text{est}}$ and the surrogate regret rate is $\psi$. Then, the target risk estimation error is of order $\psi(\epsilon_{\text{est}})$. Hence, we need to take care of both $\psi$ and $\epsilon_{\text{est}}$ to discuss the target risk estimation.
On the one hand, loss functions entailing either strongly convex or exponentially concave are known to achieve fast estimation error rates with ERM (e.g., $\mathcal{O}(1/n)$) in standard parametric setting, where the hypothesis class is of finite dimension [smooth1, mix, fast_rate, fast2]. However, poly_regret reveals that typical fast rate losses suffer at least from square-root regret rates, and thus their corresponding target risk estimation error bounds are $\mathcal{O}(1/\sqrt{n})$ at least. On the other hand, convolutional Fenchel–Young losses (which is convex smooth but not strongly convex) yield the estimation error upper bounds of order $\mathcal{O}(1/\sqrt{n})$ under the same setting, while the final target risk upper bounds are in order of $\mathcal{O}(1/\sqrt{n})$ thanks to the linear regret rate function. As a result, convolutional Fenchel–Young losses achieve target regret convergence rates that are comparable to those of existing fast-rate losses.
In more general nonparametric settings, where fast rates are often hard to achieve without additional assumptions [mendelson2008], convolutional Fenchel–Young losses achieve a target risk estimation error bound of order $\mathcal{O}(1/\sqrt{n})$. In contrast, strongly convex surrogate losses typically achieve a slower target risk estimation error rate than $\mathcal{O}(1/\sqrt{n})$ because the general nonparametric estimation error is $\mathcal{O}(1/\sqrt{n})$ but the regret rate are slower than $\psi(r)=\mathcal{O}(r)$.
While we do not intend to argue that the convolutional Fenchel–Young loss is always better, we would like to highlight that it may be a good alternative when the parametric conditions for fast rates cannot be readily justified. It also remains an open and worthwhile question whether fast rates can be obtained for our loss under additional assumptions, such as low-noise or margin conditions.
Smooth non-convex linear regret losses.
While our focus is on convex surrogates due to their favorable optimization and statistical properties, we note that certain smooth but non-convex surrogates, such as the mean absolute error (MAE) [ghosh, mao_cs] and structured comp-sum losses with MAE [mao_st], also achieve linear regret. These methods, while typically non-convex and not equipped with probability estimator, offer valuable advantages in other aspects, such as robustness to label noise, $\mathcal{H}$-consistency, and potential benefits under adversarial conditions [adv1, adv2], where non-convexity can play a meaningful role.
Appendix B Deferred Proofs in Section 3
@@
Recalling the definition of $T(\bm{p})=-\min_{t\in\widehat{\mathcal{Y}}}\left\langle\bm{p},\bm{\ell}^{\bm{\rho}}(t)\right\rangle$, which is the negative of the affinely transformed target Bayes risk in (9). It can be inferred that it is proper convex and l.s.c., and we have $\Omega_{T}=\Omega+T$.
B.1 Proof of Lemma 8
[…]
@@ B.1 Proof of Lemma 8
$T^{*}(\bm{\theta})=\sup_{\bm{p}\in\mathbb{R}^{d}}\left[\langle\bm{\theta},\bm{p}\rangle-\max_{t\in\widehat{\mathcal{Y}}}\left\langle\bm{p},-\bm{\ell}^{\bm{\rho}}(t)\right\rangle\right]=\mathbb{I}_{\mathrm{conv}\left(\{-\bm{\ell}^{\bm{\rho}}(t)\}_{t\in\widehat{\mathcal{Y}}}\right)}(\bm{\theta}),$
[…]
@@
Next we show that the infimum is indeed achieved by some $\bm{\pi}\in\Delta^{N}$. Note that for any $\bm{\theta}\in\mathbb{R}^{d}$, the set $\Theta\coloneqq\{\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi}:\bm{\pi}\in\Delta^{N}\}$ is compact and non-empty, and that $\Omega^{*}$ is l.s.c. on $\Theta$ since $\mathrm{dom}({\Omega^{*}})=\mathbb{R}^{d}$. Then the infimum of $\Omega^{*}(\bm{\theta})$ is achieved by some $\bm{\theta}^{\prime\prime}\in\Theta$, and there must exist $\bm{\pi}\in\Delta^{N}$ such that $\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi}=\bm{\theta}^{\prime\prime}$ by the definition of $\Theta$, which complete the proof of the existence of the minimizer.
Finally we see that for any $\bm{\theta}\in\mathbb{R}^{d}$, there exists $\bm{\pi}\in\Delta^{N}$ such that $\Omega_{T}^{*}(\bm{\theta})=\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi})$, which is smaller than $\infty$ since $\mathrm{dom}(\Omega^{*})=\mathbb{R}^{d}$ and $\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi}\in\mathbb{R}^{d}$. This indicates that $\mathrm{dom}(\Omega_{T}^{*})=\mathbb{R}^{d}$ because $\bm{\theta}$ is chosen arbitrarily.
B.2 Proof of Theorem 11
@@
When $\Omega$ is strictly convex, $\Omega_{T}$ is also strictly convex since $T$ is convex. Since the strict convexity holds on $\mathrm{dom}(\Omega_{T})$, $\Omega_{T}$ is further essentially strictly convex, that is, strictly convex on every convex subset of $\mathrm{dom}(\partial\Omega_{T})$ [convex, p253]. According to convex, $\Omega_{T}^{*}$ is essentially smooth, that is, differentiable throughout non-empty $\mathrm{int}(\mathrm{dom}(\Omega_{T}^{*}))$ with $\|\nabla\Omega_{T}^{*}(\bm{\theta})\|$ diverging to $\infty$ when $\bm{\theta}$ approaches a boundary point of $\mathrm{int}(\mathrm{dom}(\Omega_{T}^{*}))$, which immediately indicates the differentiability on $\mathbb{R}^{d}$.
@@
B.3 Proof of Lemma 12
Proof. According to Condition 1, the strict convexity of $\Omega$, and the proof of Theorem 11, both $\Omega_{T}^{*}$ and $\Omega^{*}$ are differentiable on $\mathbb{R}^{d}$. In addition, we have $\mathrm{dom}(\Omega)=\mathrm{dom}(\Omega_{T})$ according to (8).
@@ B.3 Proof of Lemma 12
$\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi})$ $\overset{\text{(A)}}{=}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi})^{\top}\nabla\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi})-\Omega(\nabla\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi}))$
@@
$\overset{\text{(A)}}{\leq}\bm{\theta}^{\top}\nabla\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi})+\left\langle\nabla\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi}),\mathcal{L}^{\bm{\rho}}\bm{e}_{t}\right\rangle-\Omega(\nabla\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi}))$
@@
$\overset{\text{(B)}}{=}\bm{\theta}^{\top}\nabla\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi})+\underbrace{\min_{t^{\prime}\in\widehat{\mathcal{Y}}}\langle\nabla\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi}),\bm{\ell}^{\bm{\rho}}(t^{\prime})\rangle}_{=-T\left(\nabla\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi})\right)}-\Omega(\nabla\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi}))$
$\overset{\text{(C)}}{=}\bm{\theta}^{\top}\nabla\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi})-\Omega_{T}(\nabla\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi}))$
$\overset{\text{(D)}}{\leq}\sup_{\bm{p}\in\mathrm{dom}(\Omega_{T})}\bm{\theta}^{\top}\bm{p}-\Omega_{T}(\bm{p})$
@@
where (A) owes to the optimality of $\bm{\pi}\in\Pi(\bm{\theta})$, (B) holds by the definition of $t$, and (C) holds by the definition of the convolutional negentropy $\Omega_{T}$ (8). Since $\bm{\pi}\in\Pi(\bm{\theta})$, we have $\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi})=\Omega_{T}^{*}(\bm{\theta})$ according to Lemma 8. Thus the above inequality is indeed an identity. In particular, (D) becomes an identity, which implies that the supremum of $\sup_{\bm{p}\in\mathrm{dom}(\Omega_{T})}\bm{\theta}^{\top}\bm{p}-\Omega_{T}(\bm{p})$ is achieved at $\nabla\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi})$. Since the supremum is also achieved at $\nabla\Omega_{T}^{*}(\bm{\theta})$ and the maximizer is unique according to variational and the differentiability of $\Omega^{*}_{T}$, we have
@@
Since $\bm{\pi}\in\Pi(\bm{\theta})$ and $\bm{\theta}\in\mathbb{R}^{d}$ are chosen arbitrarily, this concludes the proof.
B.4 Proof of Theorem 13
Proof. Fix any $\bm{\theta}\in\mathbb{R}^{d}$ and choose any $\bm{\pi}\in\Delta^{N}$ out of $\Pi(\bm{\theta})$. Then we have
$\geq\mathtt{Regret}_{\ell}(\varphi(\bm{\theta}),\bm{\eta})/N.$ (by definition of $\varphi$ in Definition 10)
B.5 Proof of Lemma 14